OpenStudy (anonymous):

Please help!!! quick question?? If (x + 3) / 3 = (y + 2) / 2, then x / 3 = _____.

OpenStudy (anonymous):

\[\frac{x+3}{3}=\frac{y+2}{2}\] \[\frac{x}{3}+1=\frac{y+2}{2}\] \[\frac{x}{3}=\frac{y+2}{2}-1\]

OpenStudy (anonymous):

you can also write \[\frac{x}{3}=\frac{y}{2}+1-1\] \[\frac{x}{3}=\frac{y}{2}\]

OpenStudy (ash2326):

@sani10 did you understand?

OpenStudy (anonymous):

Yes! thank you guys :)

OpenStudy (ash2326):

@satellite73 is the king here:D

OpenStudy (anonymous):

it would have been easier to write \[\frac{x+3}{3}=\frac{y+2}{2}\] \[\frac{x}{3}+1=\frac{y}{2}+1\] \[\frac{x}{3}=\frac{y}{2}\]

OpenStudy (anonymous):

king of beer.

OpenStudy (anonymous):

llite beer? (stupid joke....)